Sol Summary – November 29th>

Sol: 3

Summary Title: Winding down.

Author’s name: Enkhtuvshin “Dono” Doyodkhuu

Mission Status: Sleep schedule is slowly settling in. Still waiting for our drone to arrive. Getting used to doing EVAs. Water and power consumption needs to be closely monitored.

Sol Activity Summary: Our power went off last night due to a generator issue and we are trying our best to conserve electricity. The crew was well prepared for such occasions, and we spent the evening playing cards after dinner. Fortunately, the HAB’s propane supply is still sufficient for our meals, and we turned on the space heater in the HAB’s first floor to stay warm throughout the night. We also faced connectivity issues due to the power failure and could not send in the Greenhab report and the daily photos to Mission Support. I have just sent them in. Water consumption is also becoming a problem now, and we only have 6 days’ worth of water according to our calculations. We are now carefully monitoring our water usage.

The sol started around 8:30 AM. We had a lentil soup with vegetables for breakfast and some coffee. The first EVA started a little after 10:00 AM with Muggi as the leader with Davaa and Dono accompanying him as the designated drivers. We reached a bit farther than Cowboy Corner and returned with 48% batteries on both rovers. The rest of the day has been relaxed, and the crew is finally able to rest after a long trip and a busy initiation of the mission. Mars is starting to challenge us in a variety of ways, but the crew came prepared to survive and adapt, and morale is high.

Look Ahead Plan: We are planning to take it easy for the next few days and get accustomed to the time zone. Tomorrow, we plan to head out to Gateway to Lith while carefully monitoring the battery power of the rovers, and with constant communication with CAPCOMM. The drone should arrive sometime in the afternoon tomorrow, and we will get a better view of the environment from up above.

Anomalies in work: Power failure. Low water level. Connectivity issue. I also burnt our dinner, wasting a perfectly good black bean burger patty. Apparently it is super important to pay close attention to cooking on Mars.

Weather: Sunny with a good amount of clouds.

Crew Physical Status: 3 of the crew members are showing fatigue from the long journey. Our sleep schedule is also yet to completely settle in.

EVA: Successfully performed 1 EVA today with Muggi as the leader of the expedition.

Reports to be filed: Sol summary, Journalist report, Operations report, Greenhab report, EVA report, EVA request, Photos.

Support Requested: It has become apparent to us that the water we have left is not sufficient for the rest of the mission. We are requesting a water refill during the weekend if possible.

Copyright © The Mars Society. All rights reserved. | Main Site